The sum of the denominator and numerator of a fraction is 14. When 7 is added to each, the resulting fraction has a value of 3/4. What is the original fraction?
Answer by edjones(8007)

Let the fraction be x/y
x+y=14
(x+7)/(y+7)=3/4
4(x+7)=3(y+7)
4x+28=3y+21
4x-3y=-7
3x+3y=42
---------add
7x=35
.
x=5
y=9
.
5/9 is the original fraction.
.
Ed